Circular No. 7773

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ib IN NGC 7242 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2001ib (cf. IAUC 7768, 7772), obtained by W. Brown on Dec. 10.08 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova at maximum light. Adopting the NASA/IPAC Extragalactic Database recession velocity of 5790 km/s for the host galaxy, the supernova expansion velocity is ∼ 9700 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 1 ± 2 days after maximum light. Several titanium absorption features appear stronger than usual, indicating that this may be a subluminous event similar to SN 1999by (see Garnavich et al. 2002,

Ap.J., in press, and http://xxx.lanl.gov/abs/astro-ph/?0105490).

A. V. Filippenko and R. Chornock, University of California, Berkeley, report that inspection of CCD spectra (range 330–1000 nm), obtained on Dec. 11 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 20001ib is of type Ia. The spectral-feature age (Riess et al., op.cit.) is 5 ± 2 days past maximum brightness. The Si II 585-nm absorption might be somewhat stronger than usual, relative to the Si II 615-nm line.

COMET 51P/HARRINGTON Z. Sekanina, Jet Propulsion Laboratory, writes: “Preliminary analysis

of the reported astrometric data for the two nuclei (IAUC 7769, MPEC 2001-X45) indicates that companion D separated from primary A on Sept. 5.6 ± 3.6 TT, or 3 months after perihelion, with a relative deceleration of 59 ± 8 units of 10^{−5} the solar attraction. The two separation parameters are, however, highly correlated and probably more uncertain than their formal errors suggest. Unless too faint, the companion should have shown up in high-resolution images taken since mid-October, when it was ∼ 4′′ to the west of the primary. Predicted separation distances and position angles (0^{h} TT): 2001 Dec. 17, 12′′, 268^{o}; 2002 Jan. 6, 10′′, 263^{o}; Jan. 26, 9′′, 255^{o}; Feb. 15, 9′′, 249^{o}; Mar. 7, 11′′, 248^{o}; Mar. 27, 15′′, 251^{o}; Apr. 16, 19′′, 255^{o}; May 6, 24′′, 259^{o}.”

K. Kadota, Ageo, Saitama, Japan, reported CCD observations that indicated an intrinsic brightening of more than two magnitudes between Aug. 22 and Sept. 23 (an outburst confirmed by other observers).

2001 December 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green